OpenStudy (anonymous):

4. Is the sequence 81, 3, 1/9, … arithmetic, geometric, neither, or both? (1 point) Hint: Remember if there is a common difference between terms then it is arithmetic. If there is a common ratio between terms then it is geometric. If it doesn’t have any of those qualities then it is neither. o arithmetic o geometric o neither

hartnn (hartnn):

so, you know what is "common ratio" ??

OpenStudy (anonymous):

What is an ratio? like an fraction (1/9) ??? the word is familular

hartnn (hartnn):

yes, ratio means fraction. here, fraction with next term as numerator, and current term as denominator. and its called "common", because its common for all pair of terms.

OpenStudy (anonymous):

I put Neither because of 81 but I'm scared if its wrong, and get wrong credit on it i got a D- in math and want to bring it up.

hartnn (hartnn):

its not neither

OpenStudy (anonymous):

Aww pickle juice T_T

hartnn (hartnn):

let me give you an example: 2,4,8,16,32,.... say this is your sequence. what do we observe here ? 4/2 =2 8/4 = 2 16/8 =2 32/16 = 2 see the pattern ?? the "ratio" is "common" for all consecutive pair of terms does this makes sense ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

next one is 64?

hartnn (hartnn):

correct! ok, do you see such pattern in the sequence of your question ??

OpenStudy (anonymous):

''Common Ratio'' geometric.. O_o

hartnn (hartnn):

thats correct! :) your sequence has a common ration, and so its a geometric sequence :) any doubts ?

OpenStudy (anonymous):

ehhh but how its a ''common ratio'' if i don't see a pattern because of the 81 to the 3 how that happen?

hartnn (hartnn):

divide! is 3 divided by 81 equals 1/9 divides by 3 ??? if yes, then you have a ratio, which is common!

OpenStudy (anonymous):

am I doing it right? I got 0.03703703703704…

hartnn (hartnn):

yeah, but is it same for both ratio ?

OpenStudy (anonymous):

pooh :D 1 divided by 9 is 0.1111111… divided by 3 is 0.3703703703704…

OpenStudy (anonymous):

thnxxxx

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Can you Help with another question?

hartnn (hartnn):

sure

OpenStudy (anonymous):

19. During exercise, the recommended maximum heart rate in beats per minute is modeled by the formula, M = 176 – 0.8A, where M is the maximum heart rate and A is the person’s age. Solve the formula for A. At what age would you have a recommended maximum heart rate in beats per minute of 140? (1 point) Hint: Undo the operations to isolate A just like when we solve equations. Then substitute the 140 in for M. • 40 years • 50 years • 45 years

OpenStudy (anonymous):

I was thinking it was 40

hartnn (hartnn):

A = (176-M)/0.8

hartnn (hartnn):

A = 36/0.8 = 45 actually

OpenStudy (anonymous):

Thank you

hartnn (hartnn):

welcome ^_^